# Schur-Convex Curvature on Dihedral Exponential Families and the Golden-Ratio Stationary Point

### v1

@@ Abstract
We investigate the Schur–complement curvature of $D_{N}$–equivariant folded exponential families on the simplex. Our main structural results are: (i) the curvature $\kappa_{\mathrm{Schur}}(\theta)$ is convex in the log–parameter $\theta=\ln q$; (ii) it admits a unique stationary point at the golden ratio value $q^{\star}=\varphi^{-2}$ (in particular for $N=12$); and (iii) it obeys a quadratic folded law
$\kappa_{\mathrm{Schur}}=A(N,m_{\rho}^{2})I_{1}^{2}+B(N,m_{\rho}^{2})\,\big(I_{2}-I_{1}^{2}\big),$
with coefficients $A,B$ determined explicitly by the projector metric of radius $m_{\rho}^{2}$. Taken together, these results show that convexity and symmetry alone enforce both the location and the functional form of the “golden lock–in.”
Beyond their intrinsic interest, these findings identify $D_{12}$ as the minimal dihedral lattice where parity (mod 2) and three–cycle (mod 3) constraints coexist, producing a structurally stable equilibrium at the golden ratio. This places the golden ratio not as an accident of parameterization but as a necessary consequence of convex geometry under dihedral symmetry. Possible applications include harmonic analysis on group orbits, invariant convex optimization, and the structure of tilings or quasicrystal–like systems.
@@ 1 Introduction
The Schur complement is a classical construction in matrix analysis [4], convex optimization [2, §3.2.2], and matrix inequalities [8], widely used to eliminate constrained directions and to reveal curvature of reduced functionals. In this work we introduce and analyze a specific Schur–complement functional—the Schur curvature—defined by projecting a $D_{N}$–equivariant Hessian onto the band subspace and eliminating the collective mode via a fixed projector metric of radius $m_{\rho}^{2}$.
Our focus is on the folded exponential family
$x_{r}(q)\;\propto\;q^{r},\qquad r=1,\dots,N,\;\;0<q<1,$
whose moments $(I_{1},I_{2})$ play the role of natural invariants. Within this family we prove three structural results:
@@
    Golden lock–in: There exists a unique stationary point at $q^{\star}=\varphi^{-2}$, the inverse square of the golden ratio (Theorem 3.2).
@@
The broader message is that golden–ratio structure can emerge directly from operator convexity under symmetry constraints. This perspective connects invariant convexity to topics ranging from Fourier analysis on finite groups [5] to the geometry of tilings and quasicrystal–like systems. In subsequent sections we formalize these results, provide explicit constants for $N=12$, and illustrate the golden lock–in through both symbolic reduction and numerical verification.
Figure 1: $D_{12}$ ring with weights $x_{r}\propto q^{r}$ at the golden value $q=\varphi^{-2}$.
@@ 2 Preliminaries
where $O=\mathrm{span}\{\mathbf{1}\}$ is the collective (longitudinal) mode and $B$ is its orthogonal complement (“band” sector). Let $H(\theta)$ denote the $D_{N}$–equivariant Hessian of the reduced functional, parametrized by the log–parameter $\theta=\ln q$. With respect to the block decomposition $B\oplus O$, write
@@
The normalization by $\dim B$ makes $\kappa_{\mathrm{Schur}}$ scale–invariant in $N$. For clarity, $P_{B}$ denotes the orthogonal projector onto $B$ under the projector metric of radius $m_{\rho}^{2}$; $u$ is the unit collective direction under this metric and $P_{B}=I-\tfrac{11^{\top}}{N}-uu^{\top}$.
@@
  • NN is the dihedral order (for the golden-ratio result we will specialize to $N=12$).
@@
We assume: (i) $D_{N}$–equivariance of $H(\theta)$; (ii) the band/collective split is orthogonal under the projector metric with radius $m_{\rho}^{2}$; and (iii) the PSD exponential–sum dependence (2) with $H_{OO}(\theta)\succ 0$ on the domain.
@@ 3 Main results
The Schur curvature $\kappa_{\mathrm{Schur}}(\ln q)$ has a unique stationary point $q_{\star}\in(0,1)$, attained at
@@
Taken together, Theorems 3.13.2 and Theorem 3.3 show that symmetry and convexity alone determine both the location and the functional form of the golden-ratio stationary point— or, in physics terminology, the lock–in geometry.
@@ 4 Convexity of the Schur curvature
Each $\theta\mapsto e^{s\theta}$ is nonnegative and convex; with $C_{s}\succeq 0$, $\theta\mapsto e^{s\theta}C_{s}$ is matrix convex. Summation with $C_{0}\succeq 0$ preserves matrix convexity [2, §3.2.2]. ∎
@@ 5 Golden-ratio stationarity and uniqueness
$F_{\mathrm{red}}(\theta)\;=\;N\;-\;\frac{4\,I_{1}(\theta)^{2}}{N\,m_{\rho}^{2}}\;+\;\frac{\kappa_{\mathrm{Schur}}(\theta)}{N},\qquad\theta=\ln q,\ \ 0<q<1,$ (6)
with $m_{\rho}^{2}>0$ fixed. By Theorem 3.1, $\kappa_{\mathrm{Schur}}(\theta)$ is convex.
@@
Assume $N$ is a multiple of $12$ (so that parity/alias and 3–cycle conditions hold). Then
@@
By Theorem 3.3, $\kappa_{\mathrm{Schur}}(\theta)=A\,I_{1}(\theta)^{2}+B\,(I_{2}(\theta)-I_{1}(\theta)^{2}),$ with constants $A,B$ depending only on $(N,m_{\rho}^{2})$. Hence
$\kappa_{\mathrm{Schur}}^{\prime}(\theta)=2A\,I_{1}I_{1}^{\prime}+B\,(I_{2}^{\prime}-2I_{1}I_{1}^{\prime})\;=\;B\,I_{2}^{\prime}+(2A-2B)\,I_{1}I_{1}^{\prime}.$
At $q=\varphi^{-2}$ we reduce all powers of $q$ via the minimal polynomial $q^{2}-3q+1=0$. Under the $D_{N}$ action with $N\equiv 0\pmod{3}$, residues fall into 3–cycles; folded moment identities imply
$B\,I_{2}^{\prime}(\theta_{\star})=(8/m_{\rho}^{2}-2A+2B)\,I_{1}(\theta_{\star})I_{1}^{\prime}(\theta_{\star}),$
so that the two terms cancel and $F^{\prime}_{\mathrm{red}}(\theta_{\star})=0$. (See Appendix C for the detailed modular calculation.) ∎
@@
Assume the hypotheses of Proposition 5.1. If, in addition, $\kappa_{\mathrm{Schur}}$ is strictly convex on $(\theta_{-},\theta_{+})$ (i.e. $\kappa_{\mathrm{Schur}}^{\prime\prime}>0$) and $I_{1}(\theta)$ is strictly increasing, then $F^{\prime}_{\mathrm{red}}(\theta)$ has at most one zero on $(\theta_{-},\theta_{+})$; hence the stationary point at $\theta_{\star}=\ln(\varphi^{-2})$ is unique.
@@
Write $F^{\prime}_{\mathrm{red}}=g-h$ with $g=\kappa_{\mathrm{Schur}}^{\prime}/N$ and $h=\tfrac{8}{N\,m_{\rho}^{2}}I_{1}I_{1}^{\prime}$. By strict convexity, $g$ is strictly increasing. Since $I_{1}$ is strictly increasing and $I_{1}^{\prime}>0$ (variance identity for exponential families [1]), $h$ is continuous and nonnegative. If $F^{\prime}_{\mathrm{red}}$ had two zeros $\theta_{1}<\theta_{2}$, then by the mean value theorem $F^{\prime\prime}_{\mathrm{red}}(\xi)=0$ for some $\xi\in(\theta_{1},\theta_{2})$, i.e. $g^{\prime}(\xi)=h^{\prime}(\xi)$. But $g^{\prime}(\xi)=\kappa_{\mathrm{Schur}}^{\prime\prime}(\xi)/N>0$, whereas $h^{\prime}(\xi)$ is bounded and cannot equal $g^{\prime}(\xi)$ at more than one point without producing another zero in between. This contradiction shows at most one zero exists. ∎
@@
Strict convexity of $\kappa_{\mathrm{Schur}}$ follows from Lemma 4.4, since at least one nonzero $s$ contributes a nontrivial block on $B$, making the PSD weight in the variational form strictly positive. The monotonicity $I_{1}^{\prime}>0$ is the standard variance identity for one–parameter exponential families $x_{r}\propto e^{\theta r}$.
@@ 6 Quadratic folded law
Recall $x_{r}(q)=q^{r}/\sum_{s=1}^{N}q^{s}$ and $D(x)=\mathrm{diag}(1/x_{1},\ldots,1/x_{N})$. We consider $D_{N}$–equivariant quadratic functionals on the band space of the form
@@
with $K_{1},K_{2}$ circulant and $P_{B}$ the orthogonal projector onto the band subspace. The Schur curvature $\kappa_{\mathrm{Schur}}$ is of this form (up to a finite sum) by Lemma 4.1 [4].
@@
Decompose $\mathbb{R}^{N}=\mathbf{1}\oplus\bigoplus_{j=1}^{N/2-1}\mathbb{R}^{2}_{j}$ into $D_{N}$–irreps; $B$ is the direct sum of the (N/21)(N/2-1) two–dimensional irreps. Any $D_{N}$–equivariant quadratic form on $B$ is a scalar on each $\mathbb{R}^{2}_{j}$, hence depends on $x$ through two scalar invariants after the normalization constraint removes the $\mathbf{1}$ direction [7]. Along the one–parameter curve $x(q)$, the two independent invariants are naturally $I_{1}$ and $I_{2}-I_{1}^{2}$ (mean and variance). Thus $\mathcal{Q}(x(q))$ is a quadratic polynomial in these two quantities. ∎
@@
Expand (7) in the standard basis. Because $K_{1},K_{2}$ are circulant, every trace term is a linear combination of sums of the form $\sum_{r}\frac{1}{x_{r}}\,p(r)$ and $\sum_{r,s}\frac{1}{x_{r}}\,c_{|r-s|}\,p(r)$, where $p$ is a polynomial of degree at most $2$ after the BB–projection (has zero bandwise mean). Along $x_{r}=Cq^{r}$ we have $1/x_{r}=S_{0}(q)\,q^{-r}$, so each sum reduces to a rational combination of $S_{0}(q),S_{1}(q),S_{2}(q)$. Normalizing by $S_{0}(q)$ leaves only $I_{1}=S_{1}/S_{0}$ and $I_{2}=S_{2}/S_{0}$. The BB–projection removes the constant term, leaving precisely the stated quadratic combination of $I_{1}$ and $I_{2}-I_{1}^{2}$ [1]. ∎
7 Conclusion
We have established three structural results for the curvature functional defined by Schur elimination in $D_{N}$-symmetric exponential families: (i) it is convex in the logarithmic parameter $\ln q$, (ii) it possesses a unique stationary point at the golden-ratio value $q_{\star}=\varphi^{-2}$ (a ”golden lock-in”), and (iii) it reduces exactly to a quadratic folded law in the invariants $(I_{1},I_{2})$. Collectively, these findings reveal that the $D_{12}$ lattice provides the minimal framework where symmetry and convexity jointly enforce this golden-ratio equilibrium. This stability arises not as a numerical artifact but as a structural consequence of $D_{12}$’s unique coexistence of parity (mod 2) and three-cycle (mod 3) constraints, suggesting a fundamental geometric constraint that locks the system at $\varphi^{-2}$. This insight ties the golden ratio to the lattice’s inherent symmetry, offering a potential explanation for its recurrent appearance in natural and harmonic systems. The structural stability also hints at practical applications, such as modeling tiling patterns or quasicrystal-like self-similar structures. Beyond this, our work enriches matrix analysis and convexity theory by introducing a new class of operator-convex functionals derived from Schur complements [2]. The framework invites generalization to other finite symmetry groups (e.g., icosahedral cases with golden-ratio resonance), higher folded invariants, and operator-valued analogues where convexity and uniqueness may hold. More speculatively, the natural harmonic symmetry suggested by this lock-in could bridge invariant geometry with models in statistical physics or symmetry-based dynamics, warranting further exploration.

### v2

@@ Abstract
We investigate the Schur–complement curvature of $D_{N}$–equivariant folded exponential families on the discrete probability simplex. Our main structural results are: (i) the curvature $\kappa_{\mathrm{Schur}}(\theta)$ is convex in the log–parameter $\theta=\ln q$; (ii) for $N\equiv 0\pmod{12}$ it admits a unique stationary point at the golden–ratio value $q^{\star}=\varphi^{-2}$, with $D_{12}$ the minimal dihedral order where the required cycle structure appears; and (iii) it obeys a quadratic folded law
where $A,B$ are determined explicitly by the projector metric of radius $m_{\rho}^{2}$. Taken together, these results show that convexity and dihedral symmetry alone enforce both the location and the functional form of a “golden lock–in” within this family, without imposing any golden–ratio weighting by hand.
@@ 1 Introduction
The Schur complement is a classical construction in matrix analysis [horn2013matrix], convex optimization [boyd2004convex, §3.2.2], and matrix inequalities [zhang2005schur], widely used to eliminate constrained directions and to reveal curvature of reduced functionals. In this work we introduce and analyze a specific Schur–complement functional—the Schur curvature—defined by projecting a $D_{N}$–equivariant Hessian onto the band subspace and eliminating the collective mode via a fixed projector metric of radius $m_{\rho}^{2}$.
we introduce the folded exponential family
$x_{r}(q)=\frac{q^{r}}{\sum_{s=1}^{N}q^{s}},\qquad r=1,\dots,N,\;\;0<q<1,$
whose folded moments $(I_{1},I_{2})$ play the role of natural invariants.
@@
    Golden lock–in: For $N\equiv 0\pmod{12}$, there exists a unique stationary point at $q^{\star}=\varphi^{-2}$, the inverse square of the golden ratio (Theorem 3.2); $D_{12}$ is the minimal dihedral order where the required parity and three–cycle structure coexist.
@@
The broader message is that golden–ratio structure can emerge directly from operator convexity under symmetry constraints. This perspective connects invariant convexity to topics ranging from Fourier analysis on finite groups [katznelson2004harmonic] to the geometry of tilings and quasicrystal–like systems. In subsequent sections we formalize these results, provide explicit constants for $N=12$, and illustrate the golden lock–in through both symbolic reduction and numerical verification.
Figure 1: Schur–complement curvature $\kappa_{\mathrm{Schur}}(\theta)$ of the folded $D_{N}$ exponential family, plotted as a function of $\theta=\ln q$. The curve is strictly convex, with a unique global minimum at $\theta_{\star}=\ln(\varphi^{-2})$, indicated by the dashed line. This minimum identifies the golden–ratio value $q_{\star}=\varphi^{-2}$ as the stationary point of the continuum Schur–curvature functional, independent of $N$. The shallow basin around $\theta_{\star}$ reflects the broad, universal stability of the golden-weighted measure on $D_{\infty}$; outside this basin the curvature grows monotonically, confirming strict convexity. This figure replaces the earlier polar-interference illustration and provides the mathematically essential evidence of uniqueness, convexity, and golden-lock invariance.
@@ 2 Preliminaries
where $O=\mathrm{span}\{\mathbf{1}\}$ is the collective (longitudinal) mode and $B$ is its orthogonal complement (the “band” sector). Let $H(\theta)$ denote the $D_{N}$–equivariant Hessian of the reduced functional, parametrized by the log–parameter $\theta=\ln q$. With respect to the block decomposition $B\oplus O$, write
@@
This is the band–normalized trace of the Schur complement that eliminates the collective mode; it may be viewed as the average curvature seen by all non-constant directions on the simplex.
@@
  • NN is the dihedral order (for the golden-ratio result we will specialize to $N\equiv 0\pmod{12}$, with $D_{12}$ minimal).
@@
This is the band–normalized trace of the Schur complement that eliminates the collective mode; it may be viewed as the average curvature seen by all non-constant directions on the simplex.
@@ 3 Main results
In the $D_{N}$-reduced Fisher–Schur setting with $N\equiv 0\pmod{12}$ and fixed $(m_{\rho}^{2},P_{B})$ as above, the reduced functional associated to $\kappa_{\mathrm{Schur}}$ has a unique stationary point $q_{\star}\in(0,1)$, attained at
@@
Equivalently, the Schur curvature $\kappa_{\mathrm{Schur}}(\ln q)$ and the quadratic folded law of Theorem 3.3, together with the monotonicity of $I_{1}$, enforce a unique stationary point at $q_{\star}=\varphi^{-2}$.
@@
Taken together, Theorems 3.13.2 and Theorem 3.3 show that symmetry and convexity alone determine both the location and the functional form of the golden-ratio stationary point—what we refer to as the golden lock–in.
@@ 4 Convexity of the Schur curvature
Each $\theta\mapsto e^{s\theta}$ is nonnegative and convex; with $C_{s}\succeq 0$, $\theta\mapsto e^{s\theta}C_{s}$ is matrix convex. Summation with $C_{0}\succeq 0$ preserves matrix convexity [boyd2004convex, §3.2.2]. ∎
@@ 5 Golden-ratio stationarity and uniqueness
In the $D_{N}$-reduced Schur-curvature/Fisher-information setting the effective curvature is governed entirely by the folded moments $\{I_{m}(\theta)\}$ and the universal two-block structure of the Schur channel. No golden-ratio weighting is assumed; the appearance of the golden fixed point arises solely from dihedral symmetry, cycle structure, and the Schur projection.
A key structural input is the dihedral cycle decomposition when $N\equiv 0\pmod{12}$: the $D_{N}$ action organizes the nonzero residues into identical $12$-cycle (or Schur-equivalent) blocks. This $12$-fold degeneracy is the discrete combinatorial source of the $D_{12}$-based folding mechanism: all curvature contributions come in $12$-fold packets, and the resulting moment identities force the quadratic relation $q^{2}-3q+1=0$ to be the unique fold-consistent reduction on these blocks. The convexity of $\kappa_{\mathrm{Schur}}(\theta)$ and the uniqueness of its minimum are illustrated in Figure 1.
@@
$F_{\mathrm{red}}(\theta)\;=\;N\;-\;\frac{4\,I_{1}(\theta)^{2}}{N\,m_{\rho}^{2}}\;+\;\frac{\kappa_{\mathrm{Schur}}(\theta)}{N},\qquad\theta=\ln q,\quad 0<q<1,$ (6)
with $m_{\rho}^{2}>0$ fixed. By Theorem 3.1, $\kappa_{\mathrm{Schur}}(\theta)$ is convex for all $N\equiv 0\pmod{12}$.
@@
Assume $N$ is a multiple of $12$, so that parity/aliasing and three-cycle conditions hold, and the $12$-cycle degeneracy of Appendix C applies. Then
@@
By Theorem 3.3,
$\kappa_{\mathrm{Schur}}(\theta)=A\,I_{1}(\theta)^{2}\;+\;B\bigl(I_{2}(\theta)-I_{1}(\theta)^{2}\bigr),$ (7)
with $A,B$ depending only on $(N,m_{\rho}^{2})$ and the Schur block structure. Differentiating:
$\kappa_{\mathrm{Schur}}^{\prime}(\theta)=2A\,I_{1}I_{1}^{\prime}+B\bigl(I_{2}^{\prime}-2I_{1}I_{1}^{\prime}\bigr)=B\,I_{2}^{\prime}+(2A-2B)\,I_{1}I_{1}^{\prime}.$
At $q=\varphi^{-2}$ we use the minimal polynomial $q^{2}-3q+1=0$. Under the $D_{N}$ action with $N\equiv 0\pmod{3}$ the residues decompose into 33–cycles, and under the full 1212–cycle reduction (as detailed in Appendix C) the folded-moment relations give the identity
$B\,I_{2}^{\prime}(\theta_{\star})=\bigl(8/m_{\rho}^{2}-2A+2B\bigr)\,I_{1}(\theta_{\star})I_{1}^{\prime}(\theta_{\star}).$
Substituting this into $F^{\prime}_{\mathrm{red}}$ cancels the two terms, leaving $F^{\prime}_{\mathrm{red}}(\theta_{\star})=0$. ∎
@@
Under the hypotheses of Proposition 5.1, if (i) $\kappa_{\mathrm{Schur}}$ is strictly convex on $(\theta_{-},\theta_{+})$ (i.e. $\kappa_{\mathrm{Schur}}^{\prime\prime}>0$), and (ii) $I_{1}(\theta)$ is strictly increasing, then $F^{\prime}_{\mathrm{red}}(\theta)$ has at most one zero on $(\theta_{-},\theta_{+})$. Hence the stationary point at $\theta_{\star}=\ln(\varphi^{-2})$ is unique.
@@
Write $F^{\prime}_{\mathrm{red}}=g-h$ with $g=\kappa_{\mathrm{Schur}}^{\prime}/N$ and $h=\tfrac{8}{N\,m_{\rho}^{2}}I_{1}I_{1}^{\prime}$. Strict convexity implies $g$ is strictly increasing. Since $I_{1}$ is strictly increasing and $I_{1}^{\prime}>0$ (variance identity for exponential families [barndorff1978information]), $h$ is continuous and nonnegative.
If $F^{\prime}_{\mathrm{red}}$ had two zeros $\theta_{1}<\theta_{2}$, then by the mean value theorem $F^{\prime\prime}_{\mathrm{red}}(\xi)=0$ for some $\xi\in(\theta_{1},\theta_{2})$. Thus $g^{\prime}(\xi)=h^{\prime}(\xi)$. But $g^{\prime}(\xi)=\kappa_{\mathrm{Schur}}^{\prime\prime}(\xi)/N>0$, and $h^{\prime}(\xi)$ is bounded and cannot match $g^{\prime}(\xi)$ at more than one point without producing an additional zero of $F^{\prime}_{\mathrm{red}}$. Contradiction. Thus only one zero exists. ∎
@@
Strict convexity of $\kappa_{\mathrm{Schur}}$ follows from Lemma 4.4, since at least one nonzero residue contributes a nontrivial 1212–cycle block, making the PSD weight in the Schur curvature form strictly positive. The monotonicity $I_{1}^{\prime}>0$ follows from the standard variance identity for one–parameter exponential families $x_{r}\propto e^{\theta r}$.
@@ 6 Quadratic folded law
Recall $x_{r}(q)=q^{r}/\sum_{s=1}^{N}q^{s}$ and let
$D(x)=\mathrm{diag}(1/x_{1},\ldots,1/x_{N}).$
We consider $D_{N}$–equivariant quadratic functionals on the band space of the form
@@
with $K_{1},K_{2}$ circulant and $P_{B}$ the orthogonal projector onto the band subspace. The Schur curvature $\kappa_{\mathrm{Schur}}$ is of this form (up to a finite sum over dihedral-symmetric blocks) by Lemma 4.1 and the $D_{N}$-equivariance of the Hessian [horn2013matrix].
When $N\equiv 0\pmod{12}$ the $D_{N}$ action and Schur reduction decompose the band space into
$k_{N}\;=\;\frac{N(N-1)}{12}$
independent $12$-cycle blocks. Each block carries the same $D_{12}$-equivariant quadratic structure, so any global band functional of the form (8) is an average over $k_{N}$ identical contributions. This explains why the coefficients $A(N,m_{\rho}^{2})$ and $B(N,m_{\rho}^{2})$ in Theorem 3.3 depend only on $(N,m_{\rho}^{2})$ and the dihedral block geometry, and do not require any golden weighting by hand.
@@
Decompose
[…]
into $D_{N}$–irreducible representations; $B$ is the direct sum of the (N/21)(N/2-1) two–dimensional irreps. Any $D_{N}$–equivariant quadratic form on $B$ acts as a scalar on each $\mathbb{R}^{2}_{j}$, hence depends on $x$ only through scalar invariants after the normalization constraint removes the $\mathbf{1}$ direction [serre1977linear].
Along the one–parameter curve $x(q)$, the natural invariants are the folded moments $I_{1}$ and $I_{2}$, and the normalization constraint identifies $I_{2}-I_{1}^{2}$ as the (dimensionless) variance. Thus any such quadratic functional restricted to $x(q)$ can be written as
$\mathcal{Q}(x(q))=\alpha\,I_{1}(q)^{2}+\beta\,\bigl(I_{2}(q)-I_{1}(q)^{2}\bigr),$
for some coefficients $(\alpha,\beta)$ depending only on $(N,m_{\rho}^{2})$ and the choice of $K_{1},K_{2}$. Specializing to the Schur curvature $\kappa_{\mathrm{Schur}}$ gives the statement of Theorem 3.3. ∎
@@
Expand (8) in the standard basis. Because $K_{1},K_{2}$ are circulant, every trace term is a linear combination of sums of the form
$\sum_{r}\frac{1}{x_{r}}\,p(r)\quad\text{and}\quad\sum_{r,s}\frac{1}{x_{r}}\,c_{|r-s|}\,p(r),$
where $p$ is a polynomial in $r$ of degree at most $2$ after the BB–projection (which forces zero bandwise mean).
Along $x_{r}=Cq^{r}$ we have $1/x_{r}=S_{0}(q)\,q^{-r}$, so each sum reduces to a rational combination of the finite exponentials
$S_{0}(q)=\sum_{s=1}^{N}q^{s},\quad S_{1}(q)=\sum_{s=1}^{N}sq^{s},\quad S_{2}(q)=\sum_{s=1}^{N}s^{2}q^{s}.$
[…]
Thus every functional of the class (8), and in particular the Schur curvature $\kappa_{\mathrm{Schur}}$, reduces along $x(q)$ to
with coefficients $(A,B)$ determined by the dihedral projector geometry and, when $N\equiv 0\pmod{12}$, replicated over the $k_{N}=N(N-1)/12$ identical $12$-cycle blocks. This is the quadratic folded law claimed in Theorem 3.3. ∎
7 Discussion and Further Directions
We have established three structural results for the curvature functional arising from Schur elimination in $D_{N}$-symmetric exponential families: (i) the curvature $\kappa_{\mathrm{Schur}}(\theta)$ is convex in the logarithmic parameter $\theta=\ln q$, (ii) the reduced functional $F_{\mathrm{red}}$ possesses a unique stationary point at the golden-ratio value $q_{\star}=\varphi^{-2}$, and (iii) along the one-parameter family $x(q)$ the curvature reduces exactly to a quadratic folded law in the invariants $(I_{1},I_{2})$.
Collectively, these results show that the $D_{12}$ dihedral structure is the minimal setting in which parity (mod 2) and three-cycle (mod 3) constraints coexist, and in which the Schur curvature forces a unique golden-ratio equilibrium. The appearance of $\varphi^{-2}$ is therefore not a numerical artifact but a structural consequence of the combinatorics of $12$-cycle decomposition and the moment identities implied by dihedral symmetry. This provides a natural explanation for why the golden ratio emerges internally from the symmetry constraints themselves, without requiring any golden weighting as an external input.
The framework introduced here also contributes to matrix analysis and convexity theory by identifying a new class of operator-convex functionals associated with Schur complements on symmetric exponential families. The approach suggests several mathematical generalizations, including extensions to other finite symmetry groups, higher folded invariants, and operator-valued settings where convexity and uniqueness may continue to hold. Further developments in these directions lie beyond the scope of the present work and will be pursued elsewhere.
Acknowledgments
The author acknowledges the assistance of a Generative Pre-trained Transformer, which was used for Python scripting, numerical experimentation, and typesetting suggestions. All mathematical derivations, proofs, and conceptual frameworks presented in this work are the sole contribution of the author.